OpenStudy (anonymous):

Consider the functions f(x) = 4x2 and g(x) = 5(x + 7). Find f[g(−5)]. Please show all of your work for full credit.

OpenStudy (anonymous):

lets start with the inside function g(x)

OpenStudy (anonymous):

\[g(-5)=?\]

OpenStudy (anonymous):

10

OpenStudy (zzr0ck3r):

f(10) = ?

OpenStudy (anonymous):

40^2= 1600

OpenStudy (zzr0ck3r):

pemdas exponent first as there is no parentheses

OpenStudy (zzr0ck3r):

4*(10^2) is what it would look like

OpenStudy (anonymous):

4*(100)

OpenStudy (zzr0ck3r):

yes

OpenStudy (anonymous):

Choose the best description of how to calculate f[g(3)] if f(x) = 3x − 2 and g(x) = x2 − 1

OpenStudy (anonymous):

start with the inner function by computing g(3) then the result f(result)=3(result)-2

OpenStudy (anonymous):

s(2)-1 + 5

OpenStudy (anonymous):

=5

OpenStudy (anonymous):

Choose the correct procedure for calculating (f + g)(3) if f(x) = 3x − 2 and g(x) = x2 − 1.

OpenStudy (anonymous):

\[(f+g)(3)=f(3)+g(3)\]

OpenStudy (anonymous):

Consider the functions f(x) = 13x2 and g(x) = 7x − 5. What is the value of (f − g)(3)? 117 101 16 −101